OpenStudy (anonymous):

evaluate the power fraction f(x)=x^5/4 = 81

OpenStudy (anonymous):

Is the power of x equal to 5/4?

OpenStudy (anonymous):

It is f(x)=x^5/4 x=81

OpenStudy (anonymous):

If x=81 then subs it....ie 81^5/4...fourth root of 81^5 in other words

OpenStudy (anonymous):

If f(x) = 81 then x=81^4/5

OpenStudy (anonymous):

depends on which is the correct question

OpenStudy (anonymous):

4sq 81^5

OpenStudy (anonymous):

is that right that is what I have

OpenStudy (anonymous):

The power of x in the original quetion is 5 then?

OpenStudy (anonymous):

do you know how to do sq. root 3y+2=4 yeah I got that

OpenStudy (anonymous):

it says solve there are no examples in my book or the dvd's i have

OpenStudy (anonymous):

The question posted can be solved but please tell me what the power of x is in the question

OpenStudy (anonymous):

power of x was 5

OpenStudy (anonymous):

power of x was 5/4

OpenStudy (anonymous):

ok...then if x^5/4 =81...then x= 81^4/5

OpenStudy (anonymous):

so we make it a reciprocal

OpenStudy (anonymous):

The power 5/4 means the 5th power and then take the 4th root of the answer

OpenStudy (anonymous):

4sq.81^5

OpenStudy (anonymous):

So to get rid of it to leave x...you do the opposite ......5th root and 4th power

OpenStudy (anonymous):

5sq.81^4

OpenStudy (anonymous):

5th rt of 81^4.....only if f(x)=81 was your question

OpenStudy (anonymous):

f(x)=x^5/4 x=81

OpenStudy (anonymous):

If x=81..then simple...fourth rt of 81^5

OpenStudy (anonymous):

Good there then...sorry it took a while to sort out the question

OpenStudy (anonymous):

okay got it